# BubbleONet: A Physics-Informed Neural Operator for High-Frequency Bubble Dynamics

### v1

@@ 4.1 Branch net with adaptive activation function to address spectral bias
As discussed in the previous section, due to the large magnitude of the pressure amplitude, using $\Delta P$ as the input to the branch network is inadequate for representing the global input function space. Instead, BubbleONet directly uses the pressure profile as the input to the branch network, with the output of latent variable $B$ that includes the extracted features. As for the trunk network, it primarily encodes the time domain and the initial radius into a latent representation $T$ in a higher-dimensional space. A standard FFNN is sufficient to extract the necessary features in this context. The final output $R(P,t,R_{0})$ is computed through a linear operation on the latent variable:
@@
where $B\in\mathbb{R}^{m\times d}$ and $T\in\mathbb{R}^{n\times d}$, specifically, $n$ is the length of time, $m$ is batch size, and $d$ is the length of latent variable. To ensure the positivity of the predicted radius, a SmoothReLU (softplus) activation is applied to $R$, thus preventing nonphysical negative radius predictions.
4.2 Loss Function
@@
With the inclusion of physical constraints, the loss function consists of two components: a data loss ($\mathcal{L}_{\text{data}}$) term from the neural network predictions and an ODE loss ($\mathcal{L}_{\text{ode}}$) term derived from the governing equation. Let $\mathcal{L}$ denote the mean squared error (MSE) loss function. The total loss is defined as:
Loss $=w_{\text{data}}\mathcal{L}_{\text{data}}+w_{\text{ode}}\mathcal{L}_{\text{ode}}$ (9)
$=w_{\text{data}}\times\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\|R_{i}-\bar{R}_{i}\|_{2}^{2}+w_{\text{ode}}\times\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\|\mathcal{G}(R_{i})\|_{2}^{2},$
@@
where $\mathcal{L}_{\text{data}}$ denotes the discrepancy between the predicted radius profile $R$ and the ground truth $\bar{R}$, with $R_{i}$ representing the $i$th sample in the dataset of size $N_{d}$ and $N_{c}$ represents the number of collocation points. The symbol $\|\cdot\|_{2}$ indicates the L2 norm. The ODE loss $\mathcal{L}_{\text{ode}}$ represents the discrepancy in the governing dynamics, where $\mathcal{G}(\cdot)$ is the Runge–Kutta-based numerical operator for the governing ODE with discrete scheme form as:
$\mathcal{G}(R_{N}):=y_{N+1}-(y_{N}+dt\sum_{j=1}^{s}b_{j}k_{j}),$ (10)
where $dt$ is the time step size, $s$ is the number of Runge–Kutta stages, and $b_{j}$ is the stage weights of stage $j$. The subscript $N$ refers to the current time step in matrix form. The state matrix $y$ is defined as:
$y=\begin{bmatrix}R\\ \dot{R}\end{bmatrix}$ (11)
The computation of the intermediate stages $k_{j}$, as well as the coefficients $b_{j}$, follows the RK5(4)7M formulation and can be found in [55], where stage number is 7 with 5th-order accuracy. The weights $w_{\text{data}}$ and $w_{\text{ode}}$ are hyperparameters that balance the contributions of the data-driven loss and the physics-informed residual loss.
@@ 5 Results and Discussion
In this section, three simulation cases are discussed. The first case involves bubbles solved using R-P equation with a single initial radius of $R_{0}=50~\mu\text{m}$, representing the simplest scenario. The second case considers bubbles solved using K-M equation with the same initial radius, $R_{0}=50~\mu\text{m}$, introducing liquid compressibility effects compared to the R–P model. The third case extends the K–M model to multiple initial radii, aiming to evaluate the model’s capability in capturing bubble dynamics in higher-dimensional parameter spaces. Model training was conducted on an NVIDIA GeForce RTX 3070 GPU with 8 GB of memory, paired with an Intel Core i7-12700 CPU and 25.4 GB of system RAM. The Adam optimizer was chose for progressing the training. The hyperparameters for each of the above cases are listed in Table 2
@@ 5.1 Bubble Dynamics Governed by the R-P Equation with Single Initial Radius
The Rayleigh–Plesset equation is a nonlinear ODE that describes the dynamics of a spherical bubble in an infinite, incompressible, and viscous liquid. It serves as a fundamental model in bubble dynamics and cavitation physics [5, 6, 7]. In this section, we evaluate the model’s learning capability and performance on this simplest scenario.
@@
5.1.1 Performance
The model is trained using data simulated according to the design of experiments described in Table 1. Out of 3000 datasets, 2400 are randomly selected for training and 600 for validation, with each simulation lasting $50~\mu\text{s}$ and cut into data shape of 2000. The model is trained for 500,000 epochs. As this case involves single-radius prediction, the initial condition is not explicitly enforced. Therefore, the loss function simplifies to:
[…]
@@ 5.1.1 Performance
The weights of 1 and 100 are chosen based on the initial loss at epoch 0 to ensure that each term is reduced simultaneously. The training loss is shown in Figure 4(a), where the neural network loss $\mathcal{L}_{data}$ decreases rapidly at the beginning and gradually continues to decline as training progresses, while the governing equation loss $\mathcal{L}_{\text{ode}}$ steadily decreases throughout.
A high-frequency validation case simulated with a frequency of $1900~\text{kHz}$ and an amplitude of $2\times 10^{5}~\text{Pa}$ is used to evaluate the model performance at different training epochs, as illustrated in Figure 5. This figure provides a clear visualization of how the model gradually learns the features during training. For each validation, three plots are presented: (1) a comparison between the ground truth and the online prediction, (2) the corresponding absolute error (ground truth minus prediction), and (3) response in the frequency domain obtained via Fast Fourier Transform (FFT). In bubble dynamics, two peaks are typically observed in the frequency domain: the first corresponds to the natural frequency, and the second arises from the driving frequency of the external pressure field $P_{\infty}$. Resonance occurs when these two frequencies coincide.
At epoch 100,000, the model successfully captures the natural frequency but struggles to learn the driving frequency, resulting in an absolute error within the range of $0.05,0.05]\,\mu\text{m}$, which corresponds to approximately $0.01\%$ of the initial radius ($50\,\mu\text{m}$). At this stage, the second FFT peak is noticeably shifted from that of the ground truth, as indicated by the green circle. By epoch 300,000, the model begins to capture the driving frequency, and the absolute error near the start of the simulated time domain decreases. The second FFT peak also moves closer to the correct driving frequency. Finally, at epoch 500,000, the second FFT peak aligns well with the driving frequency, and the absolute error is further reduced to $0.02,0.02]\,\mu\text{m}$. The predicted bubble radius closely matches the ground truth. This progressive learning is also reflected in the decreasing $L_{2}$ error, which drops from $4.18\times 10^{-5}$ to $2.16\times 10^{-5}$.
Figure 5: Validation of bubble dynamics simulated with a driving frequency of $1900~\text{kHz}$ and an amplitude of $2\times 10^{5}~\text{Pa}$ after 100,000, 300,000, 500,000 epochs and the corresponding data in frequency space.
Figure 6: Validation of bubble dynamics simulated with a frequency of $900~\text{kHz}$ and an amplitude of $3\times 10^{5}~\text{Pa}$ after 100,000 epochs and the corresponding data in frequency space.
For other cases, such as bubbles subjected to lower driving frequencies, the model finds it easier to learn the underlying dynamics. For instance, in the case of a bubble simulated with a frequency of $900~\text{kHz}$ and an amplitude of $3\times 10^{5}~\text{Pa}$, the model produces satisfactory results after just 100,000 training epochs. The prediction error remains within the range [0.05,0.05][-0.05,0.05], and the corresponding MSE is $2.16\times 10^{-5}$, as shown in Figure 6. Thus, we can conclude that the learning process of BubbbleONet is from low to high in frequency domain, from left to right in the time domain.
Compared with the R-P equation, the K-M equation incorporates the compressibility of the liquid, providing a more accurate description of bubble dynamics, especially for high amplitude oscillations. This section focuses on the model’s ability to learn such complex patterns.
5.2.1 Performance evaluation
A similar dataset is constructed based on the DoE shown in Table 1, with each dataset representing bubble dynamics simulated over $55~\mu\text{s}$ and comprising 2000 sampled data points. The loss function adopts the same formulation as in Eq. 12, and its convergence behavior is depicted in Figure. 4(b). The epoch is increased to 600,000, due to the complexity introduced by the compressibility of the liquid introduced in K-M equation. Both the neural network loss and the governing equation loss exhibit a learning trend comparable to that observed in the R-P case.
Three validation cases representing driving frequencies from different domains are shown in Figure. 7. The first case corresponds to a low frequency of $523~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$. In this case, noticeable discrepancies can be observed in the radius plot, with an absolute error on the order of 0.1 and an MSE of $1.59\times 10^{-3}$. The second validation case is simulated at a frequency of $980~\text{kHz}$ and an amplitude of $3\times 10^{5}~\text{Pa}$, representing a mid-range frequency in the design of experiments. The prediction aligns closely with the ground truth, with an absolute error in the range of [0.025,0.025][-0.025,0.025] and an MSE of $3.60\times 10^{-4}$.The final case corresponds to a high driving frequency of $1927~\text{kHz}$ and an amplitude of $8\times 10^{5}~\text{Pa}$. The results demonstrate strong performance in this high-frequency regime, with an absolute error between [0.05,0.05][-0.05,0.05] and an MSE of $6.21\times 10^{-4}$. Among the three cases, the model also consistently captures the damping effect due to liquid compressibility. Moreover, the driving frequency is accurately identified in the frequency domain across all scenarios.

### v3

@@ 4.1 Single-Step DeepONet
As discussed previously, the large magnitude of the pressure amplitude renders $\Delta P$ unsuitable as an input for the branch network, as it fails to adequately represent the global input function space. Instead, our model employs the full pressure profile as the branch network input, producing a latent variable $B$ that encapsulates the extracted features. The trunk network encodes the temporal domain and initial radius into a latent representation $T$ in a higher-dimensional space, for which a standard FFNN suffices. The predicted bubble radius $R(P,t,R_{0})$ is obtained via a linear operation on the latent variables:
@@
where $B\in\mathbb{R}^{m\times d}$ and $T\in\mathbb{R}^{n\times d}$, with $n$ denoting the temporal length, $m$ the batch size, and $d$ the latent dimension. To ensure physical consistency, a SmoothReLU (softplus) activation is applied to $R$, enforcing positive radius predictions.
@@
With physical constraints incorporated, the total loss comprises of two components: a data loss ($\mathcal{L}_{\text{data}}$) from network predictions and an ODE loss ($\mathcal{L}_{\text{ode}}$) enforcing the governing dynamics. Let $\mathcal{L}$ denote the mean squared error (MSE). The total loss is expressed as
$\mathcal{L}$ $=w_{\text{data}}\mathcal{L}_{\text{data}}+w_{\text{ode}}\mathcal{L}_{\text{ode}}$ (5)
$=w_{\text{data}}\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\|R_{i}-\bar{R}_{i}\|_{2}^{2}+w_{\text{ode}}\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\|\mathcal{G}(R_{i})\|_{2}^{2},$
@@
where $\mathcal{L}_{\text{data}}$ measures the discrepancy between the predicted and reference radius profiles, and $\mathcal{L}_{\text{ode}}$ represents the residual of the governing dynamics. Here, $\mathcal{G}(\cdot)$ is the Runge–Kutta-based numerical operator,
$\mathcal{G}(R_{N}):=y_{N+1}-\left(y_{N}+dt\sum_{j=1}^{s}b_{j}k_{j}\right),$ (6)
with $dt$ as the time step, $s$ the number of stages, and $b_{j}$ the corresponding weights. The state variable is $y=[R,~\dot{R}]^{T}.$ Intermediate stages $k_{j}$ and coefficients $b_{j}$ follow the RK5(4)7M scheme [Dormand1986], featuring seven stages and fifth-order accuracy. The weights $w_{\text{data}}$ and $w_{\text{ode}}$ balance data fidelity and physics-informed regularization.
4.2 Two-Step DeepONet with Rowdy Activations
The central concept of DeepONet lies in constructing functional basis representations via the trunk network, while the branch network learns its corresponding coefficients.
Figure 4: Two step training architecture for DeepONet.
To improve training efficiency and generalization, Lee et al. [Lee2024] introduced a two-step training strategy in which QR factorization is applied to the pretrained basis functions of the trunk network before training the branch network. Building on this framework, Peyvan et al. [Peyvan2024] further enhanced the approach by employing singular value decomposition (SVD) instead of QR decomposition and by using Rowdy activation functions. Adopting this idea, we incorporate the two-step training methodology into our model; readers are referred to [Lee2024, Peyvan2024] for additional implementation details.
The overall training workflow and network architecture are illustrated in Figure 4. In the first stage, a learnable variable $A_{m}$ is introduced to emulate the latent representation of the branch network in conventional single-step training. The trunk network is first optimized to obtain its parameters $\phi^{*}$, after which the branch network is trained using the SVD-transformed outputs of the pretrained trunk network to determine its optimal parameters $\theta^{*}$.
4.3 Two-Step DeepOKAN
Kolmogorov–Arnold Neural Networks (KANs), recently proposed by Liu et al. [liu2024kan] is build upon the classical Kolmogorov–Arnold representation theorem, which originally proposed for shallow networks with limited hidden units. However, the recently formulation of KAN extends this foundational concept to contemporary deep learning architectures by leveraging training through backpropagation similar to multi-layer perceptron (MLP). This enables the framework to scale naturally to networks with any arbitrary number of layers, $L$. In this formulation, a KAN with $L$ layers is represented as a composition of $L$ learnable basis functions:
$f(\mathbf{x})=(\Phi_{L-1}\circ\Phi_{L-2}\circ\dots\circ\Phi_{1}\circ\Phi_{0})(\mathbf{x}),$
where $\mathbf{x}\in\mathbb{R}^{d}$ and $f(\mathbf{x})$ is a multivariate continuous function. Also, the functionality of basis function $\Phi$ in each layer may be defined as follows:
$\mathbf{x}_{l+1}=\Phi_{l}\mathbf{x}_{l},$
where $\mathbf{x}_{l}\in\mathbb{R}^{n_{l}}$ and $\mathbf{x}_{l+1}\in\mathbb{R}^{n_{l+1}}$ represent the input and output of the $l$-th layer, respectively, in a KAN architecture with an arbitrary number of $L$ layers, and $n_{l}$ denotes number of neurons in the $l$-th layer. The transformation $\Phi_{l}$ associated with the $l$-th layer can then be expressed in matrix form as follows:
$\Phi_{l}=\begin{bmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\dots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\dots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\dots&\phi_{l,n_{l+1},n_{l}}(\cdot)\\ \end{bmatrix}$.
In both the original formulation of the Kolmogorov–Arnold representation theorem [liu2024kan] and its recent extensions ([10763509], [Sidharth2024ChebyshevPK], [shukla2024comprehensive]), KAN have been constructed using spline basis functions based on the original formulation of KAN as well as with various alternative formulations of the basis functions with sufficient internal parameterizations. For the spline basis function, two important hyperparameters, namely, polynomial order ($k$) and grid size ($G$), are pivotal in determining the performance and expressivity of KAN architecture. The function approximation capability of KAN is fundamentally governed by the choice of basis functions, denoted as $\Phi$. A key advantage of KAN lies in their intrinsic interpretability, an aspect largely absent in conventional MLPs. KANs are adept at capturing both compositional structures and univariate functional relationships, often outperforming MLPs in terms of predictive accuracy. Moreover, they demonstrate superior parametric efficiency, achieving comparable or better results with significantly fewer trainable parameters. Despite these strengths, several practical challenges persist. Notably, the selection of important hyperparameters, such as grid resolution and spline order in the context of a spline basis, is highly problem-specific and non-trivial, highlighting the need for further research to enable more robust and automated deployment of the KAN architecture.
There has been a steep rise in utilization of the KAN architecture by leveraging their aforementioned benefits for solving PDEs in a physics-informed manner for low-dimensional [shukla2024comprehensive] as well as for high-dimensional [menon2025anant] PDEs. In the context of operator learning, Abueidd et al. [abueidda2025deepokan] proposed DeepOKAN that utilizes a basis formulation comprising of radial basis functions (RBF) as opposed to the conventional spline basis, mainly owing to their enhanced computational efficiency and better approximation properties compared to the B-splines. They evaluated their model on several mechanics problems and showed the superior performance of DeepOKAN over the DeepONet architecture with respect to the model accuracy.
Figure 5: Two step training architecture for DeepOKAN.
In this work, we propose a two-step DeepOKAN architecture that leverages the favorable properties of KAN such as interpretability and parametric efficiency while training the trunk and branch networks independently which can potentially lead to reduction in the overall training time as showcased by the previous work [Peyvan2024]. To this end, an important contribution of our work is that we provide an alternate strategy for overcoming the spectral bias while learning the operator governing the high-frequency bubble dynamics solely through continual learning using the two-step DeepOKAN architecture as opposed to utilizing the adaptive activations [jagtap2020locally] and Rowdy activations [jagtap2022deep], which are inevitable while tackling spectral bias using MLP based architectures. While Abueidd et al. [abueidda2025deepokan] proposed vanilla DeepOKAN with RBF basis, we show the need to carefully select the basis for a KAN based architecture depending on the problem that is being addressed as opposed naively selecting a fixed basis for all the problems. Figure 5 presents the schematic of the two-step DeepOKAN, which follows the two-step DeepONet framework but replaces the MLP components with more interpretable KAN architecture. The discussions henceforth will elaborate on the training procedures that were adopted for the two-step DeepOKAN architecture independently for the trunk network and the branch network.
4.3.1 Trunk Network: Overcoming spectral bias through continual learning
The trunk network is responsible for constructing a common basis that can effectively represent the bubble dynamics across a wide range of frequencies present in the training data. Furthermore, the training of the branch network is built upon the performance of the trunk network for the two-step framework, and poor representation of the frequencies present in the training data by the trunk network can propagate forward and adversely influence the training of the branch network, thereby leading to poor generalization of the overall model during inference. Hence, an efficient training procedure for the trunk network is always beneficial and leads to better generalization of the model during inference. To this end, the MLP-based two-step DeepONet (discussed in Section 4.2) utilized Rowdy activation and trained using the following formulation of the objective function.
$\phi^{*}=\text{arg min}_{\phi}\|R-A_{m}T(\phi)^{T}\|^{2}$ (7)
Here, $\tilde{R}(\phi)=A_{m}T(\phi)^{T}$ where $A_{m}$ is a trainable tensor that acts as a proxy for the branch network in its absence while $T(\phi)$ is the prediction by the trunk network and the reconstructed output prediction $\tilde{R}(\phi)$ is obtained through the tensor product of $A_{m}$ and $T(\phi)$.
However, in the context of two-step DeepOKAN, training the trunk network by utilizing the formulation of the objective shown in Equation 7 is observed to be not very effective for overcoming the spectral bias. Continual learning is a prominent learning paradigm utilized in machine learning for sequentially training a model for new tasks while preserving the old knowledge gathered from previous tasks without forgetting. In the context of KAN, Lu et al.[liu2024kan] showcased the effectiveness of the KAN architecture over the MLP counterpart for continual learning of simple 1D functions. They showed that MLP forgets old knowledge while KAN preserves them without catastrophic forgetting for a toy problem. While this is not entirely true based on the current problem we are solving, we allude to utilizing continual learning with partial reminders for overcoming the spectral bias in the context of two-step DeepOKAN. For bubble dynamics, we sort the dataset based on their dominant frequencies in High, Medium, and Low frequency ranges. After the data segregation, the model is fed with data from specific frequency ranges sequentially at regular intervals during model training. In the current scenario, the ’partial reminders’ ensure that the model does not completely forget the old knowledge from the previous frequency ranges while learning a new window of frequency range. It is also worthwhile to note that, the two step DeepOKAN solely utilizes the richness of the basis functions and doesn’t utilize any additional activation functions or learnable parameters outside of the KAN architecture for overcoming the spectral bias.
$\phi^{*}=\arg\min_{\phi}\Sigma_{i}\|R^{i}-A_{m}^{i}T(\phi)^{T}\|^{2}$ (8)
Here, $i=\{\text{High, Medium, Low}\}$ and $T(\phi)$ is prediction by the trunk network. Also note that, if $A_{m}\in\mathbb{R}^{p\times q}$ and $a+b+c=p$, then $\text{Concatenate}\{A_{m}^{\text{High}}\in\mathbb{R}^{a\times q},A_{m}^{\text{Medium}}\in\mathbb{R}^{b\times q},A_{m}^{\text{Low}}\in\mathbb{R}^{c\times q}\}=A_{m}\in\mathbb{R}^{p\times q}$. Also, $\text{Concatenate}\{R^{\text{High}}\in\mathbb{R}^{a\times q\times r},R^{\text{Medium}}\in\mathbb{R}^{b\times q\times r},R^{\text{Low}}\in\mathbb{R}^{c\times q\times r}\}=R\in\mathbb{R}^{p\times q\times r}$. The procedure for continual learning may be summarized as follows.
Learning High Frequencies:
$\phi^{*}=\arg\min_{\phi}\|R^{\text{High}}-A_{m}^{\text{High}}T(\phi)^{T}\|^{2}$ (9)
Learning Medium Frequencies while preserving the knowledge from High Frequencies:
$\phi^{*}=\arg\min_{\phi}\|R^{\text{Medium}}-A_{m}^{\text{Medium}}T(\phi)^{T}\|^{2}+\lambda_{\text{High}}\|R^{\text{High}}-A_{m}^{\text{High}}T(\phi)^{T}\|^{2}$ (10)
Learning Low Frequencies while preserving the knowledge from both Medium and High Frequencies:
$\phi^{*}=\arg\min_{\phi}\|R^{\text{Low}}-A_{m}^{\text{Low}}T(\phi)^{T}\|^{2}+\lambda_{\text{Medium}}\|R^{\text{Medium}}-A_{m}^{\text{Medium}}T(\phi)^{T}\|^{2}+\lambda_{\text{High}}\|R^{\text{High}}-A_{m}^{\text{High}}T(\phi)^{T}\|^{2}$ (11)
Here, the $\lambda$ parameter controls the ’partial reminders’ from the previous step of the continual learning that preserves knowledge gathered from the previous frequency range while learning a new frequency range. Also, it is worthwhile to note that the training for the current step is warm-started using trained parameters from the previous step. Therefore, when $\lambda=0$ the process is similar to transfer learning. However for the current problem of learning the bubble dynamics, the aim is to learn the entire frequency range present in the training data by partitioning them into manageable frequency ranges while preserving the knowledge across the entire range. Hence, $\lambda>0$ in the current work.
4.3.2 Branch Network
While the previous section discussed the training process of trunk network through continual learning, the current section discusses the details regarding training of the branch network. Here, training is performed using the full batch with the utilization of the RBF [abueidda2025deepokan] basis function.
$\theta^{*}=\arg\min_{\theta}\|B-\tilde{B}(\theta)\|^{2}$ (12)
For training the branch network, we utilize the RBF basis owing to its superior computational efficiency over the spline basis function of equivalent representational capacity. However, this was not the case for the trunk network discussed in the previous section since the RBF basis could only learn the global low-frequency structure of the bubble dynamics and failed to learn the dominant high-frequency components irrespective of the number of terms used within the RBF basis or the size of the overall KAN architecture. In a way, this emphasizes the need for careful selection of an appropriate basis function based on the underlying problem that is being solved.
@@ 5 Results and Discussion
In this section, we consider three cases: bubble dynamics governed by the R-P equation with a single initial radius $R_{0}=50~\mu\text{m}$, representing the simplest scenario; the K-M equation with a single initial radius, introducing liquid compressibility effects relative to the R-P model; and the K-M equation with multiple initial radii, aimed at evaluating the model’s capability in capturing bubble dynamics across higher-dimensional parameter spaces. We employed both single-step (Sec. 5.1 and 5.2) and two-step (Sec. 5.3) DeepONet architectures with the Rowdy activation function. Morover, the proposed two-step DeepOKAN framework (Sec. 5.4) is employed to solve the most complicated K-M equation with multiple initial radii. To emphasize its use for bubble dynamics, we refer to it as BubbleOKAN.
Model training was performed on an NVIDIA GeForce RTX 3070 GPU with 8 GB of memory, paired with an Intel Core i7-12700 CPU and 25.4 GB of system RAM. The Adam optimizer was employed to advance the training process. The hyperparameters used for each case are summarized in Table 2.
@@ 5.1 Bubble Dynamics Governed by the R-P Equation with Single Initial Radius
In this section, we evaluate the model’s learning capability and performance on this simplest scenario. The model is trained on data simulated according to the design of experiments in Table 1. Of 3000 datasets, 2400 are used for training and 600 for validation. Each simulation runs for $50~\mu\text{s}$ and is segmented into samples of size 2000. Training proceeds for 500,000 epochs. Since this case involves single-radius prediction, the initial condition is not explicitly enforced, reducing the loss function to
$\mathcal{L}=\mathcal{L}_{\text{data}}+100\,\mathcal{L}_{\text{ode}}.$ (13)
The weights (1 and 100) are chosen based on the initial epoch losses to balance the reduction of both terms.
Figure 6: Validation of bubble dynamics simulated with a driving frequency of $1800~\text{KHz}$ and an amplitude of $4\times 10^{5}~\text{Pa}$ after 100,000, 300,000, 500,000 epochs and the corresponding data in frequency space.
A high-frequency validation case with a driving frequency of $1800~\text{kHz}$ and an amplitude of $4\times 10^{5}~\text{Pa}$ is used to assess model performance at different training epochs (Figure 6). The figure illustrates how the model progressively learns key features during training. Each validation includes three plots: (1) comparison between the ground truth and the online prediction, (2) the corresponding absolute error, and (3) the frequency-domain response obtained via Fast Fourier Transform (FFT). In bubble dynamics, two peaks typically appear in the frequency spectrum: the first corresponds to the natural frequency, while the second arises from the driving frequency of the external pressure field $P_{\infty}$. At epoch 100,000, the model accurately captures the natural frequency but underestimates the driving frequency, yielding an absolute error within $0,0.1]\,\mu\text{m}$, approximately $0.2\%$ of the initial radius ($50\,\mu\text{m}$). The second FFT peak is shifted relative to the ground truth. By epoch 300,000, the model begins to learn the driving frequency, reducing the early-time absolute error and moving the second FFT peak closer to the correct position. At epoch 500,000, the second FFT peak aligns well with the driving frequency, and the absolute error decreases further to below $0.05\,\mu\text{m}$, indicating that the predicted bubble radius closely matches the ground truth.
Compared with the R–P equation, the K–M equation incorporates liquid compressibility, providing a more accurate description of bubble dynamics, particularly under high-amplitude oscillations. This section focuses on the model’s ability to learn such complex behavior.
A similar dataset is constructed based on the DoE shown in Table 1, with each dataset representing bubble dynamics simulated over $55~\mu\text{s}$ and comprising 2000 sampled data points. The loss function follows the same formulation as in Eq. 13. The number of training epochs is increased to 600,000 to account for the additional complexity introduced by liquid compressibility in the K–M equation. Both the neural network loss and the governing equation loss exhibit learning trends comparable to those observed in the R–P case.
Three validation cases representing driving frequencies from different ranges are presented in Figure 7. The first case corresponds to a low frequency of $523~\text{kHz}$ and an amplitude of $9\times 10^{5}~\text{Pa}$. Noticeable discrepancies are observed in the radius profile, with a maximum absolute error of approximately $0.2~\mu\text{m}$. The second validation case, simulated at a frequency of $980~\text{kHz}$ and an amplitude of $3\times 10^{5}~\text{Pa}$, represents a mid-range frequency in the design of experiments. Here, the prediction aligns closely with the ground truth, with a maximum absolute error of $0.04~\mu\text{m}$. The final case, corresponding to a high driving frequency of $1927~\text{kHz}$ and an amplitude of $8\times 10^{5}~\text{Pa}$, demonstrates strong performance in this high-frequency regime, with the absolute error remaining below $0.05~\mu\text{m}$. Across all cases, the model consistently captures the damping effects arising from liquid compressibility and accurately identifies the driving frequency in the frequency domain.
@@ 5.2 Bubble Dynamics Governed by the K-M Equation with Single Initial Radius
The training loss is shown in Figure 8, where the neural network loss $\mathcal{L}_{data}$ decreases rapidly at the beginning and gradually continues to decline as training progresses, while the governing equation loss $\mathcal{L}_{\text{ode}}$ steadily decreases throughout.